OpenStudy (anonymous):

Please, help. find the steady state vector for the Matrix (1st row: 0.2, 0.3, 0.5; 2nd row: 0.3, 0.6, 0.1; 3rd row : 0.5, 0.1, 0.4)

OpenStudy (anonymous):

\[\left[\begin{matrix}0.2 & 0.3 & 0.5 \\ 0.3 & 0.6 & 0.1\\0.5 & 0.1 & 0.4\end{matrix}\right]\] I stuck at factor the characteristic equation of lamda I got \[\lambda^3 - 12\lambda^2+9\lambda+110 =0\] and \[(\lambda-10)(\lambda^2 -2\lambda-11)\] Anyone helps me check whether mine is right or wrong and if it's right, why cannot I factor to find the other eigenvalues exceed 10?

OpenStudy (phi):

you should get one of the lambdas= 1 so it looks like you have the wrong characteristic equation.

OpenStudy (anonymous):

I follow the formula to take lamda, \[\lambda^3 - (trace)\lambda^2 + (A11+A22+A33)\lambda- \det (A)\]

OpenStudy (anonymous):

my pro claims that the matrix is symmetry but I don't know whether it helps or not

OpenStudy (anonymous):

I used some online tools, but this looks like the characteristic equation: -x^3 + 6/5x^2 - 9/100x - 11/100, which factors as: http://www.wolframalpha.com/input/?i=factor%28+-x%5E3+%2B+6%2F5x%5E2+-+9%2F100x+-+11%2F100%29.

OpenStudy (anonymous):

-1/100 (-1+x) (-11-20 x+100 x^2)

OpenStudy (phi):

jump is getting a lambda=1 which is a good sign.

OpenStudy (anonymous):

I'll be back in 30 minutes sorry everybody, it's emergency situation , forgive me for that

OpenStudy (anonymous):

although the original matrix is Markov, we can take 10 times to get the new matrix at integers and find out the eigenvectors. those are same.

OpenStudy (anonymous):

Hello, Hoa - the fraction in that characteristic polynomial is unimportant, it's just a constant factored out: (-1+x) (-11-20 x+100 x^2). I don't think you can scale a matrix before taking the determinant: since det(xA-lambdaI) is not equal to det(a-lambdaI).

OpenStudy (anonymous):

so, my new one is\[\left[\begin{matrix}2 & 3 & 5 \\ 3 & 6 &1\\5 & 1 & 4\end{matrix}\right]\] and base on this I make lamda equation, eigenvalue is 1/10 of the new one

OpenStudy (anonymous):

@jumping_mouse . sure. the det (A -lamda I) != det (A) but the way to get the former is what I did.

OpenStudy (anonymous):

Okay, I see what you are up to. You have a lambda = 10, so that value is your lambda = 1, right? And the associated eigenvector is (1,1,1), right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

but the eigenvector (1,1,1) should come from the eigenvalue which is NOT 1

OpenStudy (anonymous):

@jumping_mouse take a look at http://www.youtube.com/watch?v=Sf91gDhVZWU. the pro teaches me how to get characteristic equation from the original A

OpenStudy (phi):

you should be getting x^3 - 12*x^2 + 9*x + 110=0 for your char equat.

OpenStudy (anonymous):

yes, exactly what I have

OpenStudy (phi):

for you "new" matrix. The x-10 corresponds to lambda=1 in the original problem

OpenStudy (anonymous):

yes

OpenStudy (phi):

your "new" lambdas are 10, x= -2sqrt(3) + 1, x = 2sqrt(3) + 1 the lambdas for the original are 1/10 of these

OpenStudy (anonymous):

you mean the leftover is solved by x = (-b +-sqr(b^2 -4ac)) / 2a? formula of quadratic equation?

OpenStudy (phi):

you are solving \[ (\lambda-10)(\lambda^2 -2\lambda-11) = 0 \] which means \[ \lambda-10=0 \text{ or } \lambda^2 -2\lambda-11=0\]

OpenStudy (anonymous):

I know the steady state comes from the eigenvalue 1. but how to solve for that?

OpenStudy (anonymous):

as you prove, the other eigenvalue do not give out the steady state because they are !=1

OpenStudy (phi):

use the eigenvalue e.g. 10 and subtract it from the main diagonal of the "new" matrix now find the null space of the resultant matrix for other eigenvalues, do the same thing: subtract 2 sqrt(3)+1 from the main diagonal, and use elimination to find the null space. This will be the eigenvector.

OpenStudy (anonymous):

let me try and tell you my result. Can I reduced to row echelon form to get easier

OpenStudy (phi):

Can I reduced to row echelon form to get easier. Yes that is exactly what you do

OpenStudy (anonymous):

I did not have the result as my pro's one. I get vector steady is (409,25,23) while his is (1/3,1/3,1/3) :(

OpenStudy (phi):

did you row reduce -8 3 5 3 -4 1 5 1 -6 ?

OpenStudy (anonymous):

R2+R3+R1 -->R1 get 0, 0, 0 R2 stay there 3 , -4, 1 2R2 -R3 get R3 1, -9, 8

OpenStudy (anonymous):

R1 replace R3 1, -9, 8 1/23(-3 R3 + R2) 0, 1, 1 0, 0 0

OpenStudy (anonymous):

oh, oh, I got it now. thanks a TON Phi

OpenStudy (anonymous):

no, just 1,1,1 not 1/3

OpenStudy (phi):

yes, the eigenvector can be scaled by anything. so 1/3 1/3 1/3 is the same as 1 1 1 for the steady state of your problem you need more information, some initial condition (such as the states sum to 1) to pick out 1/3 1/3 1/3

OpenStudy (anonymous):

oh, yes, I know what it means. since we have eigenvector is 1, 1, 1 the steady state is (1/ 1+1+1, 1/1+1+1 and 1/1+1+1) it means )1/3,1/3,1/3) my god. eigenvector of eigenvalue 1 gives out steady state by that formula . hiiiiiiihi! thanks a lot

OpenStudy (anonymous):

Why, I don't deserve medal , phi/

OpenStudy (anonymous):

I stuck everywhere. If I don't have you guide, how can I get that result?

OpenStudy (anonymous):

anyway, I'm happy now, very very happy:)